# The fundamental localization phases in quasiperiodic systems: A unified framework and exact results

### v4

@@ Abstract
The disordered quantum systems host three types of quantum states, the extended, localized, and critical, which bring up seven distinct fundamental phases in nature: three pure phases and four coexisting ones with mobility edges, yet a unified theory with full characterization and realization of all these phases has not been developed. Here we propose a complete and unified framework based on a spinful quasiperiodic (QP) system which realizes all the fundamental localization phases, with the exact and universal results being obtained for their characterization. First, we show that the pure phases are obtained when the chiral symmetry preserves in the proposed spinful QP model, giving a criterion for the emergence of the pure phases and otherwise the coexisting ones. Further, we uncover a novel universal mechanism for the critical states that their emergence is protected by the generalized incommensurate matrix element zeros in the spinful QP model, which considerably broadens the rigorous realizations of the exotic critical states. We then show the criteria of exact solvability for the present spinful QP system, with which we construct various exactly solvable models for all distinct localization phases. In particular, we propose two novel models, dubbed spin-selective QP lattice model and QP optical Raman lattice model, to achieve all basic types of mobility edges and all the seven fundamental phases of Anderson localization physics, respectively. The experimental scheme is proposed and studied in detail to realize these models with high feasibility. This study establishes a complete and profound theoretical framework which enables an in-depth exploration of the broad classes of all fundamental localization phenomena in QP systems, and offers key insights for constructing their exactly solvable models with experimental feasibility.
@@ I Introduction
Between the extended and localized states lie the critical states, which are delocalized in both position and momentum spaces. Critical states have recently attracted extensive research interests (Hatsugai and Kohmoto, 1990; Han et al., 1994; Tanese et al., 2014; Zhou et al., 2023; Wang et al., 2016; Liu et al., 2015; Zeng et al., 2016; Yao et al., 2019; Wang et al., 2020b; Goblot et al., 2020; Tong et al., 2021; Xiao et al., 2021; Liu et al., 2022; Li et al., 2023; Wang et al., 2022b; Gonçalves et al., 2023a; Lin et al., 2023; Yang et al., 2017; Lee et al., 2023; Dai et al., 2023; Shimasaki et al., 2024b; Li et al., 2024; Wang and Zhou, 2025; Huang et al., 2025; Chen et al., 2024; Guo, 2024; Dotti et al., 2024; Yang et al., 2024; Duncan, 2024; Yao et al., 2024; Bai and Weld, 2025) due to their unique properties, such as local scale invariance, multifractality, and critical quantum dynamics. Recent advances, particularly through Avila’s global theory (Avila, 2015; Zhou et al., 2023), refine a rigorous characterization Simon and Spencer (1989); Jitomirskaya and Marx (2012) of the critical states, showing that critical states generally arise when the hopping couplings in a QP system are non-uniform and possess incommensurately distributed zeros (IDZs) in the thermodynamic limit Liu (2024). The exact theory enables unambiguous determination of critical phases and the new MEs between critical and other states Zhou et al. (2023). Furthermore, this rigorous characterization of critical states provides a well-defined foundation for incorporating many-body interactions, thereby extending the paradigm of criticality beyond the single-particle picture. Inclusion of interactions further enriches the physics of the critical phase, with multifractality of the wave function influencing both ground state properties associated with symmetry breaking (Liu, 2024; Feigel’man et al., 2007; Burmistrov et al., 2012; Zhao et al., 2019; Sacépé et al., 2020; Gonçalves et al., 2024) and the emergence of many-body critical (MBC) phases (Wang et al., 2021b, 2020b; Yoshida et al., 2024; Ahmed and Roy, 2025; Roy et al., 2025) at infinite temperature. The MBC phase, interpolating the thermal phase and many-body localization (MBL) (Pal and Huse, 2010; Nandkishore and Huse, 2015; Schreiber et al., 2015; Bordia et al., 2016, 2017a; Lüschen et al., 2017; Bordia et al., 2017b), signifies a non-ergodic but delocalized regime that violates the eigenstate thermalization hypothesis (ETH) (D’Alessio et al., 2016; Deutsch, 1991; Rigol et al., 2008; Srednicki, 1994).
@@
The recently developed quasiperiodic mosaic models (Wang et al., 2020a; Zhou et al., 2023; Gao et al., 2024; Huang et al., 2025) and their generalizations Longhi (2023); Dai et al. (2023); Lin and Gong (2024); Guo (2024); Wei et al. (2024); Longhi (2024a) advance the rigorous study of localization physics in QP systems, but many important issues remain unexplored. The extended, localized, and critical states bring up seven fundamental localization phases in nature, including three pure phases and four coexisting ones (three doubly and one triply coexisting phases). However, thus far there was no unified quantum system which is shown to host all these phases. Especially, the exactly solvable models for the complete set of such phases are crucial for the analytical study of the generic localization physics, but are currently lacking. Moreover, there is no unified theory which exactly characterizes the phase transitions between the pure localized, extended, and critical phases, and all types of (conventional and anomalous) MEs. Besides, while Avila’s global theory enables an analytic study of the spinless models, generalization of the study to the cases beyond spinless systems is not only important in fundamental physics, but also desired in experiment. For instance, the universal mechanism for critical states in the complex systems with internal degrees of freedom, such as spins, remains unknown. On the other hand, the remarkable experimental progresses in engineering the spin and orbital degrees of freedom Zhang and Liu (2018); Liu et al. (2014); Wu et al. (2016); Wang et al. (2018); Sun et al. (2018); Song et al. (2019); Lu et al. (2020); Wang et al. (2021c); Zhang et al. (2024) have enabled the realization of complex QP systems beyond spinless toy models. These crucial fundamental issues motivate us to establish a universal theory that can generically characterize spinful QP lattice systems without relying on specific microscopic details.
@@ I.1 Summary of results
In this work, we propose the generic spin-$1/2$ QP system, with which we establish a unified theoretical framework for all fundamental phases in Anderson localization physics based on the exact and universal results obtained in this system. The proposed generic QP framework unifies not only the existing 1D spinful QP models with exact analytical solutions, but also the spinless QP models through an additional Majorana representation. Most importantly, this unified system establishes a versatile platform for constructing new exactly solvable QP models exhibiting rich phase structures, including all pure phases and all coexisting phases with MEs.
The universal results are obtained in this unified framework and can be summarized as follows. First, we show with duality transformation and renormalization group method that the pure extended, localized, or critical phases are obtained when the system preserves a chiral symmetry. Accordingly, the emergence of coexisting phases with MEs can be obtained by breaking the chiral symmetry. Second, we uncover a generalized universal mechanism for critical states in the spinful quasiperiodic system, which arise from generalized incommensurate zeros (GIZs) in matrix elements. This universal mechanism opens up a much broader way of rigorously realizing and characterizing the exotic critical states. Third, we establish the exactly solvable condition, under which the 1D spinful QP system can be transformed into effectively spinless QP models of dressed particles, and then can be exactly solved. The universal results provide a powerful guiding principle to construct exactly solvable models for novel localization physics, particularly the spin-selective QP model and QP optical Raman lattice model encompassing all basic types of MEs and all seven fundamental localization phases, respectively.
@@ I.2 Organization of the work
The remainder of the paper is organized as follows. In Sec. II, we introduce the generic framework for 1D spinful QP chains, including the basic Hamiltonian and the analytical approaches employed in this study, such as dual transformations, renormalization group methods, and Avila’s global theory. In Sec. III, we present the universal results for the spinful QP chains, as organized into three theorems shown here. These universal results serve as guiding principles for constructing multiple exactly solvable models hosting distinct classes of localization physics and anomalous MEs, including a novel model that realizes the seven fundamental phases of Anderson localization in Sec. IV. In Sec. V, we propose an experimental scheme for the realization of these new models using a generic quasiperiodic optical Raman lattice. Finally, we conclude in Sec. VI, with additional details and proofs provided in Appendix.
@@ II.1 The spinful quasiperiodic model
This spinful QP system unifies various QP models, including the existing 1D spinful models with analytical solutions, and spinless models by taking the Majorana representation (see the details in Appendix A and Appendix F). More importantly, with this unified framework we shall develop a series of new and exactly solvable models which host various novel quantum phases with or without mobility edges. Below, we outline the three main theoretical approaches employed in this study.
@@ II.2 Dual transformation
The dual transformation is a non-local operation similar to the Fourier transformation, and maps the system between real space and dual momentum space. It is defined as $c_{j,s}=(1/\sqrt{L})\sum_{n}e^{-i2\pi n\alpha j}c_{n,s}$. The extended, localized, and critical states behave differently under this transformation. For an extended state, the wave function is delocalized in real space, but localized in dual momentum space, whereas the localized state is delocalized in dual momentum space but localized in real space. These opposing features are exchanged under the dual transformation. In contrast, critical states exhibit behavior distinct from the other two; they remain delocalized in both real and dual spaces, and are invariant under the transformation. For example, when the system has a QP hopping term $(V_{j}^{\mathrm{od}}a_{j}^{\dagger}a_{j+1}+\mathrm{h.c.})$, which dominates over other hopping couplings and is invariant under the dual transformation due to the presence of incommensurately distributed zeros (IDZs) in the QP hopping modulation $V_{j}^{\rm od}$, the critical states may emerge and exhibit self-dual for having incommensurate nodes corresponding to the IDZs Zhou et al. (2023); Liu (2024). These incommensurate nodes in the delocalized wave functions remain unchanged under the dual transformation, driving the delocalized states into critical states. An intuitive picture suggests that IDZs partition the 1D system into multiple segments, where the wave function reorganizes, exhibiting a self-similar pattern characteristic of the critical states.
@@ II.3 Renormalization group
The renormalization group (RG) approach (Suslov, 1982; Thouless and Niu, 1983; Ostlund and Pandit, 1984; Niu and Nori, 1986, 1990; Szabó and Schneider, 2018; Jagannathan, 2021; Gonçalves et al., 2022, 2023b) investigates the relevance of renormalized coefficients by iterating the rational approximations of the QP parameter. Specifically, one takes a rational sequence $\alpha^{(n)}=p_{n}/q_{n}$ to define a sequence of periodic Hamiltonians $H^{(n)}$, which approach the quasiperiodic regime only at the irrational limit $\alpha=\alpha^{(\infty)}$. The RG flow to $n\rightarrow\infty$ determines the properties of the states at quasiperiodic regime. With the rational sequence, the energy dispersion of $H^{(n)}$ is a periodic function of quasi-momenta $\kappa_{x}$ and $\kappa_{y}$. To systematically investigate how the relevant coefficients flow as the system size increases, we compute the characteristic determinant $P^{(n)}=|{\cal H}^{(n)}-E|$ at a target energy $E$ for a given size $L=F_{n}$, which is given by
@@
$+\epsilon_{R}^{(n)}(E,\varphi,\kappa)+T_{R}^{(n)}(E),$ (2)
where $\epsilon_{R}^{(n)}$ represents irrelevant higher harmonic terms. The asymptotic nature of eigenstates at $n\rightarrow\infty$ is determined by comparing the relative magnitudes of renormalized hopping coefficients. For extended states, the dominant hopping is along $x$ direction, yielding $\big{|}\mu_{R}^{(n)}/t_{R}^{(n)}\big{|},\big{|}V_{R}^{(n)}/t_{R}^{(n)}\big{|}\rightarrow 0$. For localized states, the dominant hopping is along $y$ direction, leading to $\big{|}t_{R}^{(n)}/V_{R}^{(n)}\big{|},\big{|}\mu_{R}^{(n)}/V_{R}^{(n)}\big{|}\rightarrow 0$. For critical states, all three hopping coefficients are equally relevant, such that $\big{|}\mu_{R}^{(n)}/V_{R}^{(n)}\big{|},\big{|}\mu_{R}^{(n)}/t_{R}^{(n)}\big{|}\geq 1$. With the RG flow we can determine the transitions between extend, localized and critical phases, and the mobility edges as the energy dependent transition points.
@@ II.4 Avila’s global theory
When the Hamiltonian in Eq. (1) can be reduced to an effectively spinless quasiperiodic (QP) chain of dressed particles, the system can be analytically characterized using Avila’s global theory Avila (2015). This theory enables the analytical calculation of the Lyapunov exponent (LE), which is the inverse of the localization length. The method provides a rigorous characterization of extended, localized, and critical states. Specifically, the non-negative LE $\gamma(E)$, quantifies the localization properties of eigenstates with energy $E$. If $\gamma(E)>0$, the state has a finite localization length $\xi(E)=\gamma^{-1}(E)$. If $\gamma(E)=0$, the state is delocalized, with a divergent localization length, corresponding to either an extended or a critical state. Critical states arise when the transfer matrix $T_{j}$ of the delocalized states becomes singular Simon and Spencer (1989); Jitomirskaya and Marx (2012). This singularity can emerge from a universal physical mechanism that the hopping couplings have IDZs Zhou et al. (2023), or the QP potential has incommensurately distributed divergence points, which effectively divides the system into infinite segments, acting as IDZs.
Furthermore, Avila’s global theory can also be employed to derive the analytical expression for the correlation length of extended states. This is achieved by performing the transformation of the system into dual space. Then one can calculate the localization length of localized states in the dual space, which is equivalent to the correlation length of the corresponding extended states in the real space.
@@
Here, $E$ is the energy of the corresponding eigenstate, $||A||$ is the norm of the matrix A, i.e. the square root of the largest eigenvalue of $A^{\dagger}A$, ${\cal T}_{n,1}=\prod_{j=1}^{n}T_{j}$. The $\epsilon$ is the imaginary part of complexified quasiperiodic phase shift $\theta+i\epsilon$. The LE is generally challenging to compute analytically, even for the simple 2-by-2 transfer matrix $T_{j}$. Nevertheless, a key result from Avila’s global theory is that with the complexification one can analytically compute the LE for all eigenstates, as long as the LE $\gamma_{\epsilon}(E)$ is convex, continuous and piecewise linear with quantized right-derivative, given by

### v5

@@ Abstract
The disordered quantum systems host three classes of quantum states, the extended, localized, and critical, which bring up seven distinct fundamental phases in nature: three pure phases and four coexisting ones with mobility edges, yet a unified theory built on universal mechanism and full realization of all these phases has not been developed. Here we propose a unified framework based on a spinful quasiperiodic (QP) system which realizes all the fundamental localization phases, with the exact and universal results being obtained for their characterization. First, we show that the pure phases are obtained when the chiral(-like) symmetry preserves in the proposed spinful QP model, giving a criterion for emergence of the pure phases and otherwise the coexisting ones. Further, we uncover a novel mechanism for the critical states that their emergence is protected by the generalized incommensurate matrix element zeros in the spinful QP model, which considerably broadens rigorous realizations of the exotic critical states. We then show criteria of exact solvability for the present spinful QP system, with which we construct various exactly solvable models for all distinct localization phases. In particular, we propose two novel models, dubbed spin-selective QP lattice model and QP optical Raman lattice model, to achieve all basic types of mobility edges and all the seven fundamental phases of Anderson localization physics, respectively. The experimental scheme is proposed and studied in detail to realize these models with high feasibility. This study establishes a complete and profound theoretical framework which enables an in-depth exploration of the broad classes of all fundamental localization phenomena in QP systems, and offers key insights for constructing their exactly solvable models with experimental feasibility.
@@ I Introduction
Between the extended and localized states lie the critical states, which are delocalized in both position and momentum spaces. Critical states have recently attracted extensive research interests (Hatsugai and Kohmoto, 1990; Han et al., 1994; Tanese et al., 2014; Zhou et al., 2023; Wang et al., 2016; Liu et al., 2015; Zeng et al., 2016; Yao et al., 2019; Wang et al., 2020b; Goblot et al., 2020; Xiao et al., 2021; Liu et al., 2022; Li et al., 2023; Wang et al., 2022b; Gonçalves et al., 2023a; Lin et al., 2023; Lee et al., 2023; Dai et al., 2023; Shimasaki et al., 2024b; Li et al., 2024; Wang and Zhou, 2025; Huang et al., 2025; Chen et al., 2024; Guo, 2024; Dotti et al., 2024; Yang et al., 2024; Duncan, 2024; Yao et al., 2024; Bai and Weld, 2025) due to their unique properties, such as local scale invariance, multifractality, and critical quantum dynamics. Recent advances, particularly through Avila’s global theory (Avila, 2015; Zhou et al., 2023), refine a rigorous characterization Simon and Spencer (1989); Jitomirskaya and Marx (2012) of the critical states, showing that critical states generally arise when the hopping couplings in a QP system are non-uniform and possess incommensurately distributed zeros (IDZs) in the thermodynamic limit Liu (2024). The exact theory enables unambiguous determination of critical phases and the new MEs between critical and other states Zhou et al. (2023). Furthermore, this rigorous characterization of critical states provides a well-defined foundation for incorporating many-body interactions, thereby extending the paradigm of criticality beyond the single-particle picture. Inclusion of interactions further enriches the physics of the critical phase, with multifractality of the wave function influencing both ground state properties associated with symmetry breaking (Liu, 2024; Feigel’man et al., 2007; Burmistrov et al., 2012; Zhao et al., 2019; Sacépé et al., 2020; Gonçalves et al., 2024) and the emergence of many-body critical (MBC) phases (Wang et al., 2021b, 2020b) at infinite temperature. The MBC phase, interpolating the thermal phase and many-body localization (MBL) (Pal and Huse, 2010; Nandkishore and Huse, 2015; Schreiber et al., 2015; Bordia et al., 2016, 2017a; Lüschen et al., 2017; Bordia et al., 2017b), signifies a non-ergodic but delocalized regime that violates the eigenstate thermalization hypothesis (ETH) (D’Alessio et al., 2016; Deutsch, 1991; Rigol et al., 2008; Srednicki, 1994).
@@
The recently developed quasiperiodic mosaic models (Wang et al., 2020a; Zhou et al., 2023; Gao et al., 2024; Huang et al., 2025) and their generalizations Longhi (2023); Dai et al. (2023); Lin and Gong (2024); Guo (2024); Wei et al. (2024); Longhi (2024a) advance the rigorous study of localization physics in QP systems, but many important issues remain unexplored. The extended, localized, and critical states bring up seven fundamental localization phases in nature, including three pure phases and four coexisting ones (three doubly and one triply coexisting phases). However, thus far there was no unified quantum system which is shown to host all these phases. Especially, the exactly solvable models for the complete set of such phases are crucial for the analytical study of the generic localization physics, but are currently lacking. Moreover, the renormalization group theory developed in Ref. Gonçalves et al. (2023b) provides important insights into the characterization of transitions between the distinct types of localization phases and mobility edges, while a unified and rigorous framework that elucidates the underlying universal physical mechanisms for all these fundamental phases remains elusive. Besides, while Avila’s global theory enables an analytic study of the spinless models, generalization of the study to the cases beyond spinless systems, or even pointing out the condition for the spinful case can be reduced to the effective spinless model, is not only important in fundamental physics, but also desired in experiment. For instance, the generic mechanism for critical states in the complex systems with internal degrees of freedom, such as spins, remains unknown. On the other hand, the remarkable experimental progresses in engineering the spin and orbital degrees of freedom Zhang and Liu (2018); Liu et al. (2014); Wu et al. (2016); Wang et al. (2018); Sun et al. (2018); Song et al. (2019); Lu et al. (2020); Wang et al. (2021c); Zhang et al. (2024) have enabled the realization of complex QP systems beyond spinless toy models. These crucial fundamental issues motivate us to establish a generic theory that can generically characterize a broad class of spinful QP lattice systems without relying on specific microscopic details.
@@ I.1 Summary of results
In this work, we propose the generic spin-$1/2$ QP system, with which we establish a unified theoretical framework for all fundamental phases in Anderson localization physics based on the exact and universal results obtained in this system, and further predict novel new localization physics. The proposed generic QP framework unifies not only the existing 1D spinful QP models with exact analytical solutions, but also the spinless QP models through an additional Majorana representation. Most importantly, this unified system establishes a versatile platform for constructing new exactly solvable QP models exhibiting rich phase structures, including all pure phases and all coexisting phases with MEs.
The main results are obtained in this unified framework and can be summarized as follows. First, we show with duality transformation and renormalization group method that pure extended, localized, or critical phases are obtained when system preserves a chiral(-like) symmetry. Accordingly, the emergence of coexisting phases with MEs can be obtained by breaking the chiral(-like) symmetry. Second, we uncover a generalized universal mechanism for critical states in the spinful quasiperiodic system, which arise from generalized incommensurate zeros (GIZs) in matrix elements. This mechanism opens up a much broader way of rigorously realizing and characterizing the exotic critical states. Third, we establish the exactly solvable condition, under which the 1D spinful QP system can be transformed into effectively spinless QP models of dressed particles, and then can be exactly solved. These results provide a powerful guiding principle to construct exactly solvable models for novel localization physics, particularly the spin-selective QP model and QP optical Raman lattice model encompassing all basic types of MEs and all seven fundamental localization phases, respectively.
@@ I.2 Organization of the work
The remainder of the paper is organized as follows. In Sec. II, we introduce the generic framework for 1D spinful QP chains, including the basic Hamiltonian and the analytical approaches employed in this study, such as dual transformations, renormalization group methods, and Avila’s global theory. In Sec. III, we present the universal results for the spinful QP chains, as organized into three theorems shown here. These results serve as guiding principles for constructing multiple exactly solvable models hosting distinct classes of localization physics and anomalous MEs, including spin-selective QP model and QP optical Raman lattice model in Sec. IV. In Sec. V, we propose an experimental scheme for the realization based on a generic quasiperiodic optical Raman lattice. Finally, we conclude in Sec. VI, with additional details and proofs provided in Appendix.
@@ II.1 The spinful quasiperiodic model
This spinful QP system unifies various QP models, including the existing 1D spinful models with analytical solutions, and spinless models by taking the Majorana representation (see the details in Appendix A and Appendix B). More importantly, with this unified framework we shall develop a series of new and exactly solvable models which host various novel quantum phases with or without mobility edges. Below, we outline the three main theoretical approaches employed in this study.
@@ II.2 Dual transformation
The dual transformation is a non-local operation similar to the Fourier transformation, and maps the system between real space and dual momentum space. It is defined as $c_{j,s}=(1/\sqrt{L})\sum_{n}e^{-i2\pi n\alpha j}c_{n,s}$. The extended, localized, and critical states behave differently under this transformation. For an extended state, the wave function is delocalized in real space, but localized in dual momentum space, whereas the localized state is localized in real space, but delocalized in dual momentum space. These opposing features are exchanged under the dual transformation. In contrast, critical states exhibit behavior distinct from the other two; they remain delocalized in both real and dual spaces. Under the dual transformation, critical states are mapped to critical states. In particular, for the self-dual cases, they are invariant under the transformation footnote_dual . For example, when the system has a QP hopping term $(V_{j}^{\mathrm{od}}a_{j}^{\dagger}a_{j+1}+\mathrm{h.c.})$, which dominates over other hopping couplings and is invariant under the dual transformation due to the presence of incommensurately distributed zeros (IDZs) in the QP hopping modulation $V_{j}^{\rm od}$, the critical states may emerge and exhibit self-dual for having incommensurate nodes corresponding to the IDZs Zhou et al. (2023); Liu (2024). These incommensurate nodes in the delocalized wave functions remain unchanged under the dual transformation, driving the delocalized states into critical states. An intuitive picture suggests that IDZs partition the 1D system into multiple segments, where the wave function reorganizes, exhibiting a self-similar pattern and multifractality of the critical states.
@@ II.3 Renormalization group
The renormalization group (RG) approach (Suslov, 1982; Thouless and Niu, 1983; Ostlund and Pandit, 1984; Niu and Nori, 1986, 1990; Szabó and Schneider, 2018; Jagannathan, 2021; Gonçalves et al., 2022, 2023b) investigates the relevance of renormalized coefficients by iterating the rational approximations of the QP parameter. In particular, Ref. Gonçalves et al. (2023b) introduced a new RG framework formulated directly in terms of renormalized coefficients, rather than conventional scaling of wave functions, and this new approach will be adopted in the present work. Specifically, one takes a rational sequence $\alpha^{(n)}=p_{n}/q_{n}$ to define a sequence of periodic Hamiltonians $H^{(n)}$, which approach the quasiperiodic regime only at the irrational limit $\alpha=\alpha^{(\infty)}$. The RG flow to $n\rightarrow\infty$ determines the properties of the states at quasiperiodic regime. With the rational sequence, the energy dispersion of $H^{(n)}$ is a periodic function of quasi-momenta $\kappa_{x}$ and $\kappa_{y}$ (in the above 2D mapped space). To systematically investigate how the relevant coefficients flow as the system size increases, we compute the characteristic determinant $P^{(n)}=\det|{\cal H}^{(n)}-E|$ at a target energy $E$ for a given size $L=q_{n}$, which yields
@@
$+\epsilon_{R}^{(n)}(E,\kappa_{x},\kappa_{y})+T_{R}^{(n)}(E),$ (2)
where $\epsilon_{R}^{(n)}$ represents irrelevant higher harmonic terms. The asymptotic nature of eigenstates at $n\rightarrow\infty$ is determined by comparing the relative magnitudes of renormalized hopping coefficients. For extended states, the dominant hopping is along $x$ axis of the 2D mapped system, yielding $\big|\mu_{R}^{(n)}/t_{R}^{(n)}\big|,\big|V_{R}^{(n)}/t_{R}^{(n)}\big|\rightarrow 0$. For localized states, the dominant hopping is along $y$ direction, leading to $\big|t_{R}^{(n)}/V_{R}^{(n)}\big|,\big|\mu_{R}^{(n)}/V_{R}^{(n)}\big|\rightarrow 0$. For critical states, all three hopping coefficients are equally relevant, such that $\big|\mu_{R}^{(n)}/V_{R}^{(n)}\big|,\big|\mu_{R}^{(n)}/t_{R}^{(n)}\big|\geq 1$. With the RG flow we can determine the transitions between extend, localized and critical phases, and the mobility edges as the energy dependent transition points.
@@ II.4 Avila’s global theory
When the Hamiltonian in Eq. (1) can be reduced to an effective spinless QP chain of dressed particles, the system may be analytically solved using Avila’s global theory Avila (2015). The global theory enables the analytical calculation of the Lyapunov exponent (LE), which is the inverse of the localization length. The method provides a rigorous characterization of extended, localized, and critical states. Specifically, the non-negative LE $\gamma(E)$, quantifies the localization properties of eigenstates of energy $E$. If $\gamma(E)>0$, the state has a finite localization length $\xi(E)=\gamma^{-1}(E)$. For $\gamma(E)=0$, the state is delocalized, corresponding to either an extended or a critical state. Critical states arise when the transfer matrix $T_{j}$ of the delocalized states becomes singular Simon and Spencer (1989); Jitomirskaya and Marx (2012). This singularity can emerge from a universal physical mechanism that the hopping couplings have IDZs Zhou et al. (2023), or the QP potential has incommensurately distributed divergence points, which effectively divides the system into infinite segments, acting as IDZs.
Furthermore, Avila’s global theory can also be employed to derive the analytical expression for the correlation length of extended states. This is achieved by performing the dual transformation of the system into dual space. Then one can study the localization length of localized states in the dual space, which is equivalent to the correlation length of the corresponding extended states in the real space.
@@
Here, $E$ is the energy of the corresponding eigenstate, $A||$ is the norm of the matrix A, i.e. the square root of the largest eigenvalue of $A^{\dagger}A$, ${\cal T}_{n,1}=\prod_{j=1}^{n}T_{j}$. The $\epsilon$ is the imaginary part of complexified quasiperiodic phase shift $\theta+i\epsilon$. The LE is generally challenging to compute analytically, even for the simple 2-by-2 transfer matrix $T_{j}$. Nevertheless, a key result from Avila’s global theory is that with the complexification one can analytically compute the LE for all eigenstates, as long as the LE $\gamma_{\epsilon}(E)$ is convex, continuous and piecewise linear with quantized right-derivative, given by